OpenStudy (anonymous):

Divide Complex Numbers 2-3i/1+2i

OpenStudy (anonymous):

\[\frac{2-3i}{1+2i}=\frac{(2-3i)(1-2i)}{(1+2i)(1-2i)} \]

OpenStudy (turingtest):

@RaijeneM can you foil that out?

OpenStudy (anonymous):

No im all LOST

OpenStudy (anonymous):

Complex numbers don't really divide so u have to rationalize the denominator in order achieve the same goal

OpenStudy (anonymous):

do you know FOIL?